OpenStudy (anonymous):

The one-to-one functions g and h are defined as follows. g{(-9,6),(-3,1),(3,-3),(8,5)} h(x)=2x+9 Find the following: g^-1(-3) h^-1(x) (h o h^-1)(-7)

OpenStudy (anonymous):

On inverse functions, the x and y values switch. For the first, switch the values, for the second, switch the variables and solve for y.

OpenStudy (triciaal):

h(0) = 2*0 + 9 = 9 h^-1 -1 = 2x + 9 x =( -1-9) / 2 = -5

OpenStudy (anonymous):

g^-1(-3)=3 h^-1(x)=-5

OpenStudy (zzr0ck3r):

\(f^{-1}\cdot f(x) = f^{-1}(f(x)) = x, \ \forall x\in \mathbb{R}\)

OpenStudy (anonymous):

@mathmale can you help me with this one?

OpenStudy (mathmale):

The one-to-one functions g is defined as follows: g{(-9,6),(-3,1),(3,-3),(8,5)} Find the following: g^-1(-3) We are asked to find the inverse function of g(x): From g(x), we want:\[g ^{-1}(x)\]which is the symbol for "the inverse of g(x)." We don't actually have g(x), so we can't find g^(-1)(x) algebraically. However, there's an important and delightful principle that we need to apply here: If f(a) = b, we have the point (a,b). If we find the inverse function \[f ^{-1}(x)\] in the usual way, then the following will be true: \[f ^{-1}(b)=a \rightarrow(b,a).\] In other words, using a as the input to f(x) produces b; using b as the input to \[f ^{-1}(x)\] produces a. So... !! If we are given points generated by g(x): {(-9,6),(-3,1),(3,-3),(8,5)} , then, simply by reversing the order of the numerals within each set of parentheses, we can get the same number of points generated by -1 g (x) : {(6,-9), (1,-3), ( you finish this, please )}

OpenStudy (anonymous):

(-3,3),(5,8)

OpenStudy (mathmale):

Cool! You've shown you understand what's happening here!

OpenStudy (anonymous):

so after working on this the greater part of the night I have g^-1(-3)= 3 h^-1(x)= -5 (h o h^-1)(-7)=-2 can anyone tell me if I did this right?

OpenStudy (anonymous):

@ganeshie8 @whpalmer4 can either of you tell me if I did this right

OpenStudy (mathmale):

g: {(-9,6),(-3,1),(3,-3),(8,5)} does have the inverse \[g ^{-1}(x):{(6,-9), (1,3), (-3,3), (5,8)}\]

OpenStudy (mathmale):

Regarding h(x)=2x+9: Let h(x)=y=2x+9, or y=2x+9 Interchange x and y: x=2y+9 Solve this for y: 2y + 9 = x - 9 = -9 (subtract 9 from both sides to isolate 2y) ------ ---- 2y = (x-9) ----- ----- 2 2 x+9 y= ----------- 2 Replace that y with h^(-1)(x) and you'll be done; you'll have the inverse of h.

OpenStudy (mathmale):

STOP here. You're not required to evaluate that inverse at any particular x value, unless I'm sorely mistaken.

OpenStudy (mathmale):

Regarding the last part of the hw problem: Since h and its inverse are just that: inverse functions, multiplying h(x) by its inverse produces simply x. If x is 5, then the desired result is also 5. Ask more questions about this if it's not clear. I like to see what you did, and to learn what your reasoning was, before I say "right" or "wrong".